OpenStudy (mazefordays090):

how do you solve for x when 47= 3x =39

OpenStudy (anonymous):

47/3=x=39/3

OpenStudy (mazefordays090):

k thx

OpenStudy (ranga):

Is it supposed to be 47 + 3x instead of 47 = 3x? How can 3x be BOTH 47 and 39?

OpenStudy (anonymous):

x has two soltion

OpenStudy (ranga):

But the way the problem is written it says 47 = 39! solve 3x = 47 is one problem solve 3x = 39 is another problem. But there is no value of x that would make 47 = 39.

OpenStudy (anonymous):

no it can be like squar root sometimes it takes two soltiion

OpenStudy (ranga):

A quadratic equation has two different x values that will satisfy the same equation. A cubic equation will have three different x values that will satisfy the same equation. But the way this problem is written is not correct.